# prompt[Prompt][List of Prompts]

### v1

@@ 4 Additional Analysis
Gold Context Size Remains Critical as Distractor Volume Increases. To evaluate the robustness of the gold context size effect under varying degrees of context noise, we systematically increased the number of distractor documents. We leveraged our flexible adaptation of NuminaMath1.5 to run experiments with 5, 10, and 15 distractors, approximately 25k, 50k, and 75k distractor tokens, respectively.
Figure 7 shows that performance continues to be strongly influenced by gold context size, regardless of distractor volume. This reinforces that gold context size remains a dominant variable, even when noise levels change.
@@
Summary. These additional analyses confirm that the observed effects are not artifacts of a single benchmark or setup. Small gold contexts not only reduce performance, but also magnify positional bias. These effects are more severe in noisy environments and domain-specific tasks.
@@ 5 Related Work
Positional biases in LLMs. Position bias, the tendency of LLMs to over- or under-attend to different parts of the input, has emerged as a fundamental challenge. Prior work has identified several variants: primacy bias, where early content is favored  [13]; recency bias, where later content dominates  [15]; and U-shaped bias, where mid-context is under-attended  [14]. These effects persist across model architectures, alignment strategies  [14], extended context lengths  [29], and even, to some extent, in internal representations  [30]. Our work contributes to this literature by introducing a new dimension: we show that the size of the gold context modulates the strength of positional bias. Specifically, smaller gold contexts are significantly more vulnerable to primacy effects, while larger contexts confer greater robustness to positional variation.
The closest work to our setup is recent work of Levy et al.  [31] who study needle-in-a-haystack performance under variable input lengths. While both works investigate positional dynamics in noisy settings, our approach holds the distractor context fixed and instead varies the gold context size, allowing us to isolate the effects of gold signal sparsity.
@@
Frameworks for long-context evaluation. Evaluation strategies for long-context reasoning have evolved from synthetic toy tasks to richer, more realistic setups. Long-Range Arena  [11] introduced standardized tasks for comparing various transformer variants. Recent benchmarks explore broader benchmarking variations  [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], such as document synthesis  [12, 49], document-level retrieval  [21], citation verification  [50], and biomedical reasoning  [51, 6]. Most of these setups use the “needle-in-a-haystack” formulation  [52, 53] where a small relevant span must be retrieved from a large set of distractors. Some efforts push beyond this setup, incorporating aggregation, multi-hop inference  [54, 55], or mixed-modality inputs  [56]. Our work builds on this direction by adapting natural, domain-specific datasets to simulate realistic multi-agent aggregation within a ‘needle-in-a-haystack” framework due to its practical relevance.
@@
Mitigation strategies for position bias. Several mitigation approaches have been proposed to reduce position sensitivity in LLMs. These include compressing or abstracting context  [57], distilling long-context information into weights  [58], reweighting attention via calibration  [59], modifying positional encoding schemes  [60, 61], and fine-tuning on debiased distributions  [62]. While some methods mitigate positional biases, many introduce side effects  [63], leaving long-context generalization an ongoing challenge. Our contribution is diagnostic rather than corrective. We uncover a novel interaction between input structure (gold context size) and positional bias severity, showing that simply increasing the amount of gold evidence can systemically impact position bias. Whether existing mitigation strategies can address this effect remains an open question for future work.
@@ 6 Discussion, Limitations, and Conclusion
Why does gold context size strongly affect aggregation accuracy? Our findings reveal two interconnected factors: First, we hypothesize that larger gold contexts attract attention by offering a higher density of semantically relevant tokens, making them more prominent within distracting content. This richer semantic environment helps models retrieve relevant signals and reduces positional sensitivity. The effect is especially pronounced in domain-specific tasks, where coherent reasoning chains in larger contexts help models follow structured logic needed for accurate answers.
Practical implications of our findings. While prior work has studied factors like positional bias and distractor count, we highlight an overlooked and less controllable factor: gold context size. Therefore, practitioners should recognize that aggregation quality is sensitive to context length variations, even when retrieval mechanisms functions as expected. Practitioners can address this by strategically balancing retrieved document sizes and accounting for potential biases against shorter contexts.
Limitations of our study. First, we did not explicitly control the proportion of gold context within the total context window. Instead, we fixed distractor lengths to better reflect real-world conditions, resulting in varying gold-to-distractor ratios. This may confound whether performance differences stem from gold context size alone or its relative share. Second, while our benchmarks and distractors were curated for realism and domain diversity, only the CBB dataset used a real-world retriever; NQ and NM relied on synthetic setups. Future work should address these.
@@
Appendix A Extended Experimental Details
@@ A.1 Original Benchmarks
  • Source: CARDBiomedBench on Hugging Face and its BiomedSQL variant on Hugging Face. Distractor documents were retrieved using a multi-agent retrieval system at NIH, which retrieves content from: (1) Google search over NIH domains, (2) PubTator3.0  [64], (3) the Human Gene Mutation Database (HGMD)  [65], and (4) NCBI gene and variant pages  [66].
@@
  • Preprocessing: None, the distractor and gold documents are as-is from the retriever.
@@
  • Source: NQ with evidence spans aligned to Knowledge Intensive Language Tasks (KILT) on Hugging Face. Gold documents were loaded using the Ai2 ir_datasets python package  [67] and distractors were sourced from HELMET on Hugging Face.
@@
  • Preprocessing: Filtered to retain only examples with ‘complete’ and ‘verified’ fields for question, final answer, structured solution, and long-form generation. DeepSeekR1 generations were truncated to the final 5,000 tokens using GPT-4o tiktoken  [68] tokenization to normalize document length across tasks. Distractors sampling was among the other questions excluded duplicates. Sizes of gold and distractors were strung into a pseudo-document by including ‘The answer to {question} is {gold_document} to give them context.
A.2 LLM Configuration
@@
API Providers.
We used the following service providers for model access:
  • GPT models (GPT-4o, GPT-4o-mini) were accessed via the Azure OpenAI service.
  • Gemini models (Gemini-2.0-Flash, Gemini-2.0-Flash-Lite) were accessed via the Google AI GenAI SDK, using the official genai Python client.
@@
  • LLaMA model < 70B parameters (LLaMA-3.1-8B-Instruct) was evaluated locally using the meta-llama/Llama-3.1-8B-Instruct checkpoint, loaded via Hugging Face transformers. All local evaluations were conducted on the NIH High-Performance Computing (HPC) Biowulf cluster  [69], leveraging GPU nodes for inference.
Prompting and Evaluation Configuration.
Prompts were benchmark-specific and standardized across model types. All models were queried with max_tokens=256 and temperature=0.0. Provider-specific configurations (e.g., safety settings for Google GenAI, and device mapping for HuggingFace) were handled automatically during model initialization. See the code and YAML config files for full details.
Grading LLMs.
For CARDBiomedBench, an additional grading LLM was used to assess answer correctness via BioScore using GPT-4o, as done by the authors. It was instantiated using the same infrastructure and configurations as the primary LLMs, with max_tokens=10.
A.3 Metrics
@@
Quality Rate.
We evaluate responses to the CBB tasks following their proposed BioScore framework, an LLM-as-a-judge metric implemented with GPT-4o. Each response is scored on a 3-point scale according to the BioScore prompt A.4, and a score $\geq 2$ is considered factually correct. The Quality Rate is computed as the proportion of responses meeting this threshold.
@@
SubEM.
For NQ we utilized substring exact match, which assigns a score of 1.0 if any normalized ground truth string is a subspan of the model’s response (after normalization), and 0 otherwise. This is a correctness signal used by previous work on this data.
math-verify.
Evaluated with math-verify  [70], a symbolic equivalence checker that parses LaTeX boxed answers and verifies correctness through structured math expression comparison. Parsing and verification are done using an extraction and comparison pipeline derived from the Math-Verify toolkit.
Error Bars.
All plots showing aggregate scores (e.g., Figure 4) report 90% confidence intervals (CIs) estimated via non-parametric bootstrapping over tasks. Given $N$ scores, we resample with replacement 1,000 times and compute the middle 90% interval from the resulting bootstrap distribution.
A.4 Prompts.
@@
The NaturalQuestions prompt is adapted from previous work’s experimental methods  [14, 21].
@@
Appendix B Extended Results
We provide extended baselines for all benchmarks in Figure 8 and main results for CBB in Figure  9, for NQ in Figure  10, and for NM in Figure  11. Additionally, we provide positional curves for all models in Figure  12.
B.1 Baselines
(a) CARDBiomedBench
(b) NaturalQuestions
(c) NuminaMath1.5
Figure 8: Baseline performance when viewing distractors only, closed book (no documents), and varying sizes of gold. This confirms both (1) models perform poorly without the gold documents and (2) performance is near perfect when viewing any size of gold document.
B.2 CARDBiomedBench
Figure 9: CARDBiomedBench performance for each model and size of gold for varying positions in the context window (0.0, 0.25, 0.5, 0.75, 1.0), the average across all positions, and baseline performance when seeing gold only. Higher scores (light yellow) is more desirable than low scores (dark red), 90% CI are reported.
B.3 NaturalQuestions
Figure 10: NaturalQuestions performance for each model and size of gold for varying positions in the context window (0.0, 0.2, 0.4, 0.6, 0.8, 1.0), the average across all positions, and baseline performance when seeing gold only. Higher scores (light yellow) are more desirable than low scores (dark red), 90% CI are reported.
B.4 NuminaMath1.5
Figure 11: NuminaMath1.5 performance for each model and size of gold for varying positions in the context window (0.0, 0.2, 0.4, 0.6, 0.8, 1.0), the average across all positions, and baseline performance when seeing gold only. Higher scores (light yellow) are more desirable than low scores (dark red), 90% CI are reported.

### v2

@@ 4.1 Gold Context Position
4.2 Answer Token Repetition
Figure 5: Performance across all models on CARDBiomedBench when bucketing tasks into two categories: answer token repetition less than or equal to, and greater than, 1.5. Smaller golds yield lower accuracy compared to larger golds across both bins. Error bars are 95% confidence intervals, with some wider due to sample size.
If larger gold contexts contained the exact answer tokens more frequently, this could partly explain the phenomenon we observed—repeated encoding of the answer tokens would make it easier for the model to detect and attend to them. To test this, we compute answer token repetition, defined for an answer $a$ and a context $c$ as: $\text{AnsTokRepetition}(a,c)=\frac{1}{|T(a)|^{*}}\sum_{t\in T(c)}\mathbf{1}[t\in T(a)],$ where $T(x)$ are the tokens of $x$, $|T(x)|^{*}$ is the number of unique tokens in $x$, and $\mathbf{1}[\cdot]$ is the indicator function. Distributions of this metric across gold sizes and benchmarks are provided in Appendix D.2. We then bucket tasks using the median repetition value (1.5) and compare performance within each bin for CBB (Figure 5).
When answer repetition is low, larger gold contexts still outperform smaller ones, showing that repetition alone cannot explain the size effect. When repetition is high, performance improves for medium and large golds but not for small ones, indicating that repetition further amplifies the size effect. While repetition is beneficial for LLM performance, it is insufficient to explain the observed size effect. Small gold contexts remain consistently harder to detect and use.
4.3 Gold-to-distractor Ratio
Figure 6: Performance across all models on CARDBiomedBench when bucketing tasks by the mean (0.029) for gold-to-distractor ratio. Smaller golds yields lower accuracy compared to larger golds. Error bars are 95% confidence intervals, some are larger due to small sample size in that bin.
Given a fixed distractor budget, changing the size of the gold context also changes the proportion of gold tokens to distractor tokens within a context window. For a gold context $g$ and a distractor set $D$, we can compute: $\text{Gold-to-Distractor Ratio}(g,D)=\frac{|T(g)|}{\sum_{d\in D}|T(d)|},$ where $T(x)$ are the tokens of $x$ and $|T(x)|$ is the total number of tokens. This raises the question of, if the positive effect of larger gold contexts is due to the size itself, or the increased ratio? Bucketing the tasks using the mean ratio (0.029) into similar ranges of gold-to-distractor ratio, we can see if size still has an effect when the ratio is held constant. Figure 6 shows just this, and larger golds consistently outperform smaller ones within bins. Even after controlling for gold-to-distractor ratio, gold context size remains a strong indicator of performance.
4.4 Distractor Volume
To evaluate the robustness of the gold context size effect under varying degrees of context noise, we systematically increased the number of distractor documents. We leveraged our adaptation of NuminaMath1.5 to run experiments with 5, 10, and 15 distractors, approximately 25k, 50k, and 75k distractor tokens, respectively. Figure 7 shows that performance is strongly influenced by gold context size, regardless of distractor volume. This reinforces that size remains a dominant variable, even when noise levels change.
@@ 4.4 Distractor Volume
4.5 Domain Specificity of Tasks
Figure 8: Positional sensitivity for CARDBiomedBench. For each model and gold context size, we compute the performance range across positions. Smaller gold contexts exhibit much higher sensitivity (larger ranges), especially in domain-specific tasks. Larger gold contexts yield more stable performance across positions.
The effects of gold context size are notably amplified in domain-specific tasks compared to general knowledge. Figure 8 quantifies this for CBB by measuring the range in model performance across different gold context positions. (The results for other benchmarks are in Appendix C.2). For each model and gold size, we compute the performance range as the difference between maximum and minimum scores across all positions:
Range $=\mathop{\max}_{i\in\{1,\dots,n\}}\text{perf}(\text{position}_{i})$ (1)
$\;\;-\;\mathop{\min}_{i\in\{1,\dots,n\}}\text{perf}(\text{position}_{i})$
4.6 Joint Modeling of Gold-Size & Confounders with Multivariate Regression
To assess the independent contribution of gold size relative to other factors, we fit a multivariate logistic regression predicting correctness from gold size, answer-token repetition, gold-to-distractor ratio, and position across all tasks and models. Earlier sections analyzed these variables in isolation, but in realistic long-context settings they co-vary, making it unclear which effects genuinely stem from gold size. The joint regression provides a unified estimate of each factor’s partial effect.
Model specification. For each benchmark, we regress the binary correctness outcome on four categorical predictors:
$\Pr\!\big(\mathrm{correct}\big)=\sigma\!\left(\beta_{0}+\beta_{\text{size}}\,X_{\text{size}}+\beta_{\text{rep}}\,X_{\text{rep}}+\beta_{\text{ratio}}\,X_{\text{ratio}}+\beta_{\text{pos}}\,X_{\text{pos}}\right)$ (2)
All variables are encoded categorically for direct interpretability. Repetition and ratio are binarized into high vs. low relative to within-benchmark means, and all predictors enter the model through their corresponding indicator variables $X_{\text{size}},X_{\text{rep}},X_{\text{ratio}},X_{\text{pos}}$.
Here are how these variables defined: (1) Gold Size ($X_{\text{size}}$). A categorical factor encoding gold size, with small as the reference and indicators for medium and large. $\beta_{\text{size}}$ gives the change in log-odds when moving from small to larger contexts. (2) Answer Token Repetition ($X_{\text{rep}}$). Token overlap with the answer, binarized at the benchmark mean into low (reference) vs. high. $\beta_{\text{rep}}$ gives the log-odds difference between the two levels. (3) Gold-to-Distractor Ratio ($X_{\text{ratio}}$). The proportion of gold to distractor tokens in the context window, binarized at the benchmark mean into low (reference) vs. high. The coefficient $\beta_{\text{ratio}}$ quantifies the effect of moving from a lower share of gold tokens to a higher one. (4) Position ($X_{\text{pos}}$). The gold context’s normalized start position, with $0.0$ as the reference. Indicators represent other positions, and $\beta_{\text{pos}}$ captures how placement in the window affects correctness.
Predictor Coef 95% CI
Intercept 1.6459 [1.571, 1.720]
Gold Size
  md vs. sm 0.3933 [0.336, 0.451]
  lg vs. sm 0.5354 [0.466, 0.605]
Answer Token Repetition
  high vs. low 0.5924 [0.511, 0.674]
Gold-to-Distractor Ratio
  high vs. low 0.3699 [0.279, 0.461]
Position
  0.25 vs. 0.0 -0.8087 [-0.896, -0.721]
  0.50 vs. 0.0 -0.9001 [-0.987, -0.813]
  0.75 vs. 0.0 -0.8930 [-0.980, -0.806]
  1.00 vs. 0.0 -0.8701 [-0.957, -0.783]
Table 1: Logistic Regression Results (CBB)
Results. We show CBB results in Table 1 where we see that gold size has a strong independent effect: medium and large contexts significantly boost correctness after adjusting for all confounders. Repetition and ratio also help, while position has the largest negative impact—performance drops sharply when the gold appears anywhere other than at the start of the window. CBB performance improves reliably with larger gold contexts, in ways not explained by repetition, ratio, or position alone.
Appendix D.4 also reports all three benchmarks. Across all benchmarks, gold context size remains a significant, independent predictor of correctness even after controlling for repetition, ratio, and position. Medium and large golds consistently increase the log-odds of producing the correct answer relative to small ones.
@@ 5 Related Work
Long-context biases in LLMs. Position bias, the tendency of LLMs to over- or under-attend to different parts of the input, has emerged as a fundamental challenge. Prior work has identified several variants: primacy bias, where early content is favored (wang2023primacy); recency bias, where later content dominates (zheng2023large); and U-shaped bias, where mid-context is under-attended (liu2024lost). These effects persist across model architectures, alignment strategies (liu2024lost), extended context lengths (lee2024can; veselipositional), and, to some extent, in internal representations (lu2024insights). Our work contributes to this literature by introducing a new dimension: we show that the size of the gold context modulates the strength of positional bias.
A few recent works have examined different trade-offs in NIAH settings. levy2024task study variable input lengths and show that, for a fixed gold context, adding more distractors degrades performance. This is essentially the inverse of our setup, where we fix the distractor content and systematically vary the length of the gold context. The parallel work of levy2025more investigates the case where all documents have equal length while the total context length is held fixed. They find that longer documents (with fewer total documents) make the gold easier to discover. This is complementary to our work: instead of varying absolute length under equal-length constraints, we keep distractors fixed and vary the relative size of the same gold evidence. Finally, dai2024deniahlincontextfeaturesinfluence study NIAH performance in synthetic key–value retrieval tasks, focusing on varying answer-span lengths and task lengths. In contrast, we hold the question and answer fixed and vary the size of the surrounding gold context in open-ended QA, enabling a controlled analysis of how gold context size alone affects aggregation robustness.
@@
Frameworks for long-context evaluation. Long-context evaluation has progressed from synthetic toy tasks to increasingly realistic settings. Early work such as Long-Range Arena (tay2021long) introduced standardized tasks for comparing transformer variants. Subsequent benchmarks expanded this space (bai2024longbench; li2024loogle; gao2025uniah; li2025larabench; modarressi2025nolima; ling2025longreason; zhang2024toolbehonest; ye2024analobench; yen2024helmet), exploring document synthesis (shaham2022scrolls; shaham2023zeroscrolls), document-level retrieval (yen2025helmetevaluatelongcontextlanguage), citation verification (zhang2024longcite), and biomedical reasoning (adams2024longhealth; cui2025curie). Most adopt a needle-in-a-haystack (NIAH) formulation (kamradt2023needle; hsieh2024ruler), where a small relevant span must be recovered from distractors. Others move beyond strict NIAH by incorporating aggregation, multi-hop inference (zhuang2025docpuzz; katsis2025mtrag), or mixed-modality inputs (wu2025longmemeval). Our work builds on this trajectory by adapting natural, domain-specific datasets to simulate realistic NIAH aggregation.
@@
Mitigation strategies for long-context biases. Prior work proposes various methods for reducing positional sensitivity, including context compression (jiang2024longllmlingua), distilling long-context information into model weights (cao2025infiniteicl), attention calibration (hsieh2024found), modified positional encodings (zhang2024found; zheng2024dape), and debiased fine-tuning (xiong2024artificial). While these methods can alleviate positional bias, many introduce side effects (zhao2024understanding), and robust long-context generalization remains challenging. Our contribution is diagnostic rather than corrective: we uncover a novel bias connected to the gold context size. Whether existing mitigation strategies can address this effect remains an open question.
@@ 6 Discussion, Limitations, and Conclusion
Why does gold context size strongly affect the accuracy? At a high level, the finding here is pretty intuitive: Larger gold contexts spread semantically relevant information across more tokens, making the signal more resilient to positional noise and less likely to be overlooked amid distractors. However, we do not claim this is the only mechanism involved. Our findings suggest that interacting factors may be at play here, as discussed in §4—including positional dynamics, answer-context richness, evidence competition, and domain effects. While the overall trend (large > medium > small) is consistent across all analyses, the precise interaction among these factors remains open and is an important direction for future work.
Practical implications of our findings. While prior work has studied factors like positional bias and distractor count, our results highlight an overlooked factor: the effect of heterogeneity in the size of evidence documents. Aggregation quality can degrade sharply when small-but-critical evidence is mixed with much longer retrieved passages. This naturally arises in retrieval-augmented systems or agentic pipelines where an aggregator must select a single piece of evidence from a concatenation of heterogeneous sources. Practitioners may mitigate this by ensuring that the documents are balanced in lengths before being aggregated or by avoiding pipelines that combine extremely short and very long contexts, thereby reducing size-induced attention asymmetries.
Limitations of our study. We fixed distractor lengths to better reflect real-world conditions, resulting in varying gold-to-distractor ratios. This may confound whether performance differences stem from gold context size alone or its relative share. Parallel work has investigated this, offering complementary results to ours (levy2025more). Second, the tasks we use could be further grounded in real deployment scenarios. Third, our study focuses on aggregating a single needle; extending this to multi-needle or multi-hop settings is an important next step. Future work should address these.
@@
Appendix A Appendix
Appendix B Extended Experimental Details
@@ B.1 Original Benchmarks
  • Source: CARDBiomedBench on Hugging Face and its BiomedSQL variant on Hugging Face. Distractor documents were retrieved using a multi-agent retrieval system, which retrieves content from: (1) Google search over NIH domains, (2) PubTator3.0 (pubtator3), (3) the Human Gene Mutation Database (HGMD) (Stenson2020HGMD), and (4) NCBI gene and variant pages (ncbi2025).
@@
  • Preprocessing: None, the distractor and gold contexts are as-is from the retriever.
@@
  • Source: NQ with evidence spans aligned to Knowledge Intensive Language Tasks (KILT) on Hugging Face. Gold contexts were loaded using the Ai2 ir_datasets python package (macavaney:sigir2021-irds) and distractors were sourced from HELMET on Hugging Face.
@@
  • Preprocessing: Filtered to retain only examples with ‘complete’ and ‘verified’ fields for question, final answer, structured solution, and long-form generation. DeepSeekR1 generations were truncated to the final 5,000 tokens using GPT-4o tiktoken (tiktoken) tokenization to normalize document length across tasks. Distractors sampling was among the other questions and excluded duplicates. Sizes of gold and distractors were strung into a pseudo-document by including ‘The answer to {question} is {gold_document} to give them context.
B.2 Task Creation
Here we provide details on the gold context construction as discussed in §2.2. Specifically, in Figure 9 we show examples of the gold contexts, of different sizes, for each dataset.
B.3 LLM Configuration
@@
API Providers. We used the following service providers for model access:
  • GPT models (o3-mini, GPT-4o, GPT-4o-mini) were accessed via the Azure OpenAI service.
  • Gemini models (Gemini-2.5-Flash, Gemini-2.0-Flash, Gemini-2.0-Flash-Lite) were accessed via the Google AI GenAI SDK, using the official genai Python client.
  • DeepSeek-R1 and Phi-4-reasoning were accessed via the Azure AI Inference service.
@@
  • LLaMA model < 70B parameters (LLaMA-3.1-8B-Instruct) was evaluated locally using the meta-llama/Llama-3.1-8B-Instruct checkpoint, loaded via Hugging Face transformers. All local evaluations were conducted on the NIH High-Performance Computing (HPC) Biowulf cluster (biowulf), leveraging GPU nodes for inference.
Prompting and Evaluation Configuration. Prompts were benchmark-specific and standardized across model types. All non-reasoning models were queried with max_tokens=256 and temperature=0.0. Provider-specific configurations (e.g., safety settings for Google GenAI, and device mapping for HuggingFace) were handled automatically during model initialization. See the code and YAML config files for full details. Reasoning models were queried with max_tokens=2048 and their default generation params, to allow for reasoning. Reasoning models were additionally given instructions to encourage grounding their answer in the retrieved documents, to prevent relying on internal knowledge.
Grading LLMs. For CARDBiomedBench, an additional grading LLM was used to assess answer correctness via BioScore using GPT-4o, as done by the authors. It was instantiated using the same infrastructure and configurations as the primary LLMs, with max_tokens=10.
B.4 Metrics
@@
Quality Rate. We evaluate responses to the CBB tasks following their proposed BioScore framework, an LLM-as-a-judge metric implemented with GPT-4o. Each response is scored on a 3-point scale according to the BioScore prompt B.5, and a score $\geq 2$ is considered factually correct. The Quality Rate is computed as the proportion of responses meeting this threshold.
@@
SubEM. For NQ we utilized substring exact match, which assigns a score of 1.0 if any normalized ground truth string is a subspan of the model’s response (after normalization), and 0 otherwise. This is a correctness signal used by previous work on this data.
math-verify. Evaluated with math-verify (huggingface2025mathverify), a symbolic equivalence checker that parses LaTeX boxed answers and verifies correctness through structured math expression comparison. Parsing and verification are done using an extraction and comparison pipeline derived from the Math-Verify toolkit.
Error Bars. All plots showing aggregate scores (e.g., Figure 3) report 90% confidence intervals (CIs) estimated via non-parametric bootstrapping over tasks. Given $N$ scores, we resample with replacement 1,000 times and compute the middle 90% interval from the resulting bootstrap distribution.
B.5 Prompts.
@@
The NaturalQuestions prompt is adapted from previous work’s experimental methods (liu2024lost; yen2025helmetevaluatelongcontextlanguage).
@@
Appendix C Extended Results
As discussed in §2.3 we evaluated gold-only baselines where each gold context size (small, medium, large) was presented alone, without distractors. This verified that all variants were independently sufficient to solve the task and that downstream performance drops are due to aggregation effects (e.g., distractor interference or gold placement). We provide baselines for all benchmarks in LABEL:apdx:baselines Figure LABEL:fig:baseline_heatmap. Additionally, we provide full results for CBB in Figure LABEL:fig:cbb_heatmap, for NQ in Figure LABEL:fig:nq_heatmap, and for NM in Figure LABEL:fig:nm_heatmap. Finally, we provide positional curves for all models in Figure 11.
C.1 Performance by Position